OpenStudy (anonymous):

what is equal to √50a 6b7

OpenStudy (solomonzelman):

\[\sqrt{50a^6b^7}\]THIS?

OpenStudy (anonymous):

yes, thanks

OpenStudy (anonymous):

how did you get the little 6, I couldn't post that right. :(

OpenStudy (solomonzelman):

\[\sqrt{25} \times \sqrt{2} \times \sqrt{(a^3)^2} \times \sqrt{(b^3)^2} \times \sqrt{b}\]See how I am braking it?

OpenStudy (solomonzelman):

\[5 \times \sqrt{2} \times a^3 \times b^3 \times \sqrt{b}\]---->\[5a^3b^3\sqrt{2b}\]

OpenStudy (anonymous):

you are so helpful

OpenStudy (solomonzelman):

Anytime, @darlenedarlene!